# Towards the Training of Deeper Predictive Coding Neural Networks

### v2

@@ 4 Methods
In this section, we first study the phenomenon of energy imbalance across different layers, and then use the derived insights to define different formulations of time-dependent covariances that address it. In detail, we introduce two methods, called spiking precisions and decaying precisions, each offering distinct strategies to better distribute the energy across the model by dynamically updating the precisions. We also introduce a variation of the learning phase, which leverages neural activities at initialization to perform a better update of the parameters and improve overall model performance. We conclude with a novel formulation of the batch normalization, more suitable to the training of PCNs. We will show that this novel formulation will not have a large impact in terms of performance when used in combination with the standard formulation of PC. However, it will result in a large gain when tested with our newly proposed methods. To facilitate a better understanding of our proposed algorithms, Figure 1(a) presents a flowchart that intuitively illustrates the modules discussed in subsequent sections. Additionally, Figure 1(b) provides a visualization of the covariance matrices.
Figure 2: Layer-wise energy distribution and accuracy comparison between BP and PCNs in a VGG10 on the CIFAR10 dataset. Colored curves represent the total energy of the individual layers of the model (or, the squared error of every layer for BP). The vertical lines represent the train and test accuracies of the model.
To study the energy imbalance across different levels of the network, we have tracked the total energy of each layer during training, along with the test and training loss, and compared it against that of BP. We have performed a broad study that tests multiple models, datasets, and setups, which we mostly report in the supplementary material, while presenting in Figure 2 the plots of the best performing models. As BP does not have a proper definition of energy, we have used the squared error of every layer computed during the backward pass, equivalent to the error of $PC$ when it comes to the update of the weight parameters. As it can be observed, in PCNs there is a significant energy imbalance, where early layers have multiple orders of magnitude less energy than later layers. This does not happen in BP-trained models, that exhibit a more uniform energy distribution across layers, with even the layer of lowest energy above $10^{-2}$. This set of experiments shows the potential reason why deep PC models do not perform well. Our proposed methods (second row) seem to slightly mitigate such an energy imbalance, despite it being still very significant. However, this partial mitigation is significant enough to allow the model to fit the training set. In the rest of the section, we describe the methods that we have used to address this major bottleneck.
@@
Training predictive coding models involves a critical trade-off between stability and the efficient propagation of error signals. On the one hand, large learning rates $\alpha$ for the neural activities can lead to instabilities: most of the best results in the field have been obtained using a small learning rate, such as $\alpha=0.05$ (Pinchetti et al., 2024). On the other hand, such small learning rates can exponentially slow down the propagation of error signals across model layers, as noted in the supplementary material of (Song et al., 2020). To this end, we propose to modulate the precision having a spike proportional to the learning rate the first time the energy — initially concentrated in the output neurons — reaches a specific layer. In terms of temporal scheduling, this happens at $l=L-t$. For a network with $L$ layers and $T$ inference steps, the proposed spiking precision hence is:
@@
Intuitively, the spikes allow the energy to be well propagated from the last to the first layer during the first $L$ iterations, with the other updates happening as usual. While this method has the additional advantage of not introducing any new hyperparameter, it fails to address the second problem that we highlighted: that the large energy and continuous updates make the neural activities diverge too much from the forward pass.
Decaying Precision.
Building upon the idea of balancing stability and efficient error propagation, we introduce a decaying precision strategy. Here, the precision of different layers decays exponentially over time, as defined as follows:
Here, the numerator sum serves as a normalization term that ensures that the sum of the layer-wise precisions over time is equal to one, that is, $\sum_{t=1}^{T}(\Sigma_{t}^{l})^{-1}=1$. The denominator $e^{-k\cdot(l-L+t)}$ allows lower layers to receive larger weights when activated ($l\geq L-t$), thereby helping to achieve a more balanced energy distribution during the inference phase, k is a hyperparameter that controls the strength of this balancing effect, the search range is presented in Table 5. It also ensures that each layer experiences a significant boost in precision precisely when the energy from the output first reaches that layer $(l=L-t)$. When $l<L-t$, we set $\Sigma^{l}_{t}=1$.
@@
Due to large prediction errors that we find in the last layers, the neural activities observed at the end of the inference process tend to significantly deviate from their initial feed-forward values. But the feedforward values are the ones that are then used for predictions. We then conjecture that synaptic weight updates based on $\mathbf{x}^{l}_{T}$ could potentially introduce errors that accumulate with network depth, leading to performance degradation in deeper architectures. To assess whether the proposed conjecture is correct, we introduce a new method for updating the weight parameters that uses both the starting and final states of neurons, according to the energy function defined as
@@
Our method makes sure that weight adjustments stay connected to the initial feed-forward predictions while incorporating the refined representations obtained through iterative inference. This approach has the advantage of maintaining stability during learning and prevents the accumulation of errors in deeper layers, which is crucial for scaling PC networks, but also the disadvantage of storing information in memory, which is then used for the weight update only, making it slightly less bio-plausible than the original formulation. A similar energy function has been used in a different way in a previous work, where the authors used it to guide the update of the neural activities instead of the weight updates (Whittington and Bogacz, 2017). The pseudocode is provided in Algorithm 1.
Algorithm 1 Learning a dataset $\mathcal{D}=\{\mathbf{o}_{i},\mathbf{y}_{i}\}$ with forward updates
1:$\mathbf{x}^{0}_{0}$ is fixed to $\mathbf{o}_{i}$ and $\mathbf{x}^{L}_{t}$ is fixed to $\mathbf{y}_{i}$ for every mini-batch.
2:for $t=1$ to $T$ do
3:     for every level $\ell$ do
4:         Update $\mathbf{x}_{t}^{l}$ to minimize $E_{t}$.
5:     end for
6:     if $t=T$ then
7:         Update $\mathbf{W}^{l}$ to minimize $\tilde{E}_{T}$.
8:     end if
9:end for
@@
where $\gamma$ and $\beta$ are learnable parameters, and $\mu_{B}$ and $\sigma^{2}_{B}$ are the mean and variance of the minibatch $B$. At test time, it uses statistics $\mu_{r}$ and $\sigma^{2}_{r}$ learned through exponential moving averages.
Figure 3: Test accuracy between models with the standard formulation of BN, without BN, and our proposed BF.
However, when applied directly to PCNs, BatchNorm fails to yield similar improvements. We hypothesize that this failure results from the iterative inference phase, where processing the same batch multiple times leads to a possible overfitting of the layer statistics. To address this issue, we propose BatchNorm Freezing (BF), a modification that freezes the states of the BatchNorm state during the inference phase, and updates running statistics exclusively during the learning phase while still using batch statistics for normalization during inference iterations. Our experimental results demonstrate that this modification not only preserves network convergence but also enhances performance, suggesting that its ability to stabilize activation distributions effectively supports the energy-based optimization in PC architectures.
@@ 5 Experiments
In this section, we test our proposed methods combined on architectures with more than $7$ layers, and show that we are able to reach performance comparable to those of BP when trained on models with the same complexity. To provide a comprehensive evaluation, we test them on CIFAR-10/100 (Krizhevsky et al., 2009). We consider the following three baselines: standard PC (described in Section  3), PC with center nudging (PC-CN), the best performing algorithm according to previous studies (Pinchetti et al., 2024; Scellier et al., 2024), and standard BP. [^2] Here, we follow the standard PC literature, where performance are compared against BP, and differs from standard works in the Eqprop literature, that uses BPTT as a baseline. As architectures, we use VGG-like models (Simonyan and Zisserman, 2014), which are deep convolutional models followed by feedforward layers. Similarly to the setups of the aforementioned works, in this work we only consider models with a single feedforward layer after the convolutional layers. This allows us to check whether our proposed methods still suffer from performance degradation when the depth of the model increases. In all cases, we perform a large number of hyperparameter searches, report the best test accuracy obtained with early stopping, averaged over $5$ runs. All the details needed to reproduce the results can be found in the supplementary material.
Figure 4: Test accuracies of different algorithms on the CIFAR10 dataset, tested on models of different depths. From left to right, we have first models trained with BP, second models trained with PC, while the rest represents combinations of the novel proposed in this work.
Performance Degradation.
In the first experiment, we test some of the proposed algorithms on VGG models with different depths, and report the bar plots of their test accuracies as a function of the number of layers in Figure 4. The plots show that both the original formulation of PC and the one that uses center nudging significantly drop in test accuracy when the depth of the model is increased. In contrast, our proposed methods combining precision and forward updates (D+F, S+F) demonstrate excellent performance, notably avoiding the significant accuracy degradation as model depth increases. Particularly, spiking precision with forward updates is the only combination of methods that exhibits performance comparable to BP across VGG5/7/10 models. We will later see that this result will be consistent with the more complex experiments, where this method is still the one achieving the best performance overall. In models without forward update, consistent with previous findings, using center nudging enhances algorithm accuracy, with more pronounced effects as the number of model layers increases in most cases. However, we found that once forward update is added, center nudging ceases to yield performance benefits. To this end, from now on we will only report experiments that use the standard formulation of PC, without any nudging.
Test Accuracies.
@@
Table 1: Test accuracies of the different algorithms on different datasets.
Dataset Algorithm VGG5 VGG7 VGG10 VGG5BF VGG7BF VGG10BF
Decaying Precision (D) $87.91^{\pm 0.22}$ $81.14^{\pm 0.19}$ $84.87^{\pm 0.19}$ $88.55^{\pm 0.09}$ $80.00^{\pm 0.10}$ $81.68^{\pm 0.18}$
Forward Update (F) $88.79^{\pm 0.04}$ $87.43^{\pm 0.30}$ $87.33^{\pm 0.14}$ $87.34^{\pm 0.13}$ $85.91^{\pm 0.09}$ $79.43^{\pm 0.20}$
D+F $89.32^{\pm 0.14}$ $89.34^{\pm 0.09}$ $89.43^{\pm 0.18}$ $\mathbf{90.37^{\pm 0.13}}$ $\mathbf{91.48^{\pm 0.12}}$ $91.46^{\pm 0.12}$
S+F $\mathbf{89.45^{\pm 0.18}}$ $\mathbf{90.08^{\pm 0.21}}$ $92.07^{\pm 0.10}$ $89.30^{\pm 0.13}$ $90.89^{\pm 0.04}$ $\mathbf{93.27^{\pm 0.10}}$
CIFAR100 (Top-1) BP $66.28^{\pm 0.23}$ $65.36^{\pm 0.15}$ $\mathbf{69.35^{\pm 0.16}}$ $67.39^{\pm 0.25}$ $67.67^{\pm 0.11}$ $71.25^{\pm 0.21}$
Forward Update (F) $65.34^{\pm 0.07}$ $64.50^{\pm 0.14}$ $61.69^{\pm 0.79}$ $64.69^{\pm 0.17}$ $63.52^{\pm 0.32}$ $60.33^{\pm 0.30}$
D+F $66.10^{\pm 0.09}$ $64.86^{\pm 0.10}$ $66.54^{\pm 0.12}$ $\mathbf{67.56^{\pm 0.25}}$ $67.27^{\pm 0.21}$ $69.81^{\pm 0.22}$
S+F $\mathbf{66.49^{\pm 0.15}}$ $\mathbf{66.34^{\pm 0.22}}$ $63.35^{\pm 0.12}$ $67.16^{\pm 0.16}$ $\mathbf{67.71^{\pm 0.10}}$ $\mathbf{72.02^{\pm 0.12}}$
CIFAR100 (Top-5) BP $85.85^{\pm 0.27}$ $84.41^{\pm 0.26}$ $\mathbf{88.74^{\pm 0.08}}$ $89.56^{\pm 0.08}$ $\mathbf{90.05^{\pm 0.13}}$ $\mathbf{92.10^{\pm 0.12}}$
Decaying Precision (D) $81.59^{\pm 0.13}$ $74.00^{\pm 0.30}$ $83.13^{\pm 0.74}$ $88.82^{\pm 0.07}$ $78.93^{\pm 0.29}$ $81.16^{\pm 0.36}$
Forward Update (F) $85.48^{\pm 0.08}$ $84.05^{\pm 0.07}$ $76.73^{\pm 0.92}$ $88.05^{\pm 0.16}$ $87.45^{\pm 0.24}$ $83.33^{\pm 0.19}$
D+F $85.85^{\pm 0.10}$ $83.80^{\pm 0.20}$ $86.10^{\pm 0.21}$ $\mathbf{89.84^{\pm 0.17}}$ $89.74^{\pm 0.12}$ $91.24^{\pm 0.07}$
S+F $\mathbf{86.36^{\pm 0.11}}$ $\mathbf{84.53^{\pm 0.15}}$ $86.84^{\pm 0.07}$ $89.57^{\pm 0.09}$ $89.62^{\pm 0.18}$ $\mathbf{92.10^{\pm 0.10}}$

### v3

@@ 4 Methods
In this section, we first study the phenomenon of energy imbalance across different layers, and then use the derived insights to propose time-dependent covariances that address it. In detail, we propose spiking precisions, a method that better distributes the energy across the model by dynamically updating the precisions as soon as individual neurons are reached by the energy term. In practice, we show that this largely improves the performance of models trained with PC and iPC. In the case of iPC, spiking precisions allow us to reach performance comparable to that of backprop. To further boost the performance of standard PC, we introduce a variation of the weight update rule, which leverages neural activities at initialization to perform a better update of the parameters and improve overall model performance. Figure 1(a) presents a flowchart that intuitively illustrates the modules discussed in subsequent sections, while Figure 1(b) provides a visualization of the covariance matrices.
Figure 2: Normalized layer-wise energy distribution and accuracy comparison between BP and PCNs in a VGG10 on the CIFAR10 dataset. Colored curves represent the total energy of the individual layers of the model (or, the squared error of every layer for BP). The vertical lines represent the train and test accuracies of the model.
To study the energy imbalance across different levels of the network, we have tracked the normalized total energy of each layer during training, along with the test and training accuracy, and compared it against that of BP. We have performed a broad study that tests multiple models, datasets, and setups, which we mostly report in the supplementary material, while presenting in Figure 2 the plots of the best performing models. As BP does not have a proper definition of energy, we have used the squared error of every layer computed during the backward pass, equivalent to the error of $PC$ when it comes to the update of the weight parameters.
Results.
In models trained with PC, there is a significant energy imbalance, where early layers have up to $18$ orders of magnitude less energy than later layers, while in models trained with iPC this phenomenon is less pronounced but still present, as the layer with less energy has an energy of about $10^{-5}$ . This does not happen in BP-trained models, which exhibit a more uniform energy distribution across layers. In fact, the first layer has energy above $10^{-2}$ almost the whole training. This set of experiments shows the potential reason why deep PC models do not perform well. In the bottom row, we show how combinations of our proposed methods mitigate such an energy imbalance, with the best performing ones being iPC with spiking precisions, and PC with spiking precisions and a novel update rule we will discuss later. In both cases, the first layer has an energy above the layer of lowest energy above $10^{-3}$. These two methods are also the only ones reaching test accuracy slightly better than those of backprop.
4.1 Algorithmic Contributions
@@
Training predictive coding models involves a critical trade-off between stability and the efficient propagation of error signals. On the one hand, large learning rates $\alpha$ for the neural activities can lead to instabilities: most of the best results in the field have been obtained using a small learning rate, such as $\alpha=0.05$ (Pinchetti et al., 2024). On the other hand, such small learning rates can exponentially slow down the propagation of error signals across model layers, as noted in the supplementary material of (Song et al., 2020). To this end, we propose to modulate the precision having a spike proportional to the learning rate the first time the energy — initially concentrated in the output neurons — reaches a specific layer. In terms of temporal scheduling, this happens at l=Ltl=L-t. For a network with $L$ layers and $T$ inference steps, the proposed spiking precision hence is:
@@
Intuitively, the spikes allow the energy to be well propagated from the last to the first layer during the first $L$ iterations, with the other updates happening as usual.
@@
Due to large prediction errors that we find in the last layers, the neural activities observed at the end of the inference process tend to significantly deviate from their initial feed-forward values. But the feedforward values are the ones that are then used for predictions. We hence conjecture that synaptic weight updates based on $\mathbf{x}^{l}_{T}$ could potentially introduce errors that accumulate with network depth, leading to performance degradation in deeper architectures. To assess whether the proposed conjecture is correct, we introduce a new method for updating the weight parameters that uses both the starting and final states of neurons, according to the energy function defined as
@@
Our method makes sure that weight adjustments stay connected to the initial feed-forward predictions while incorporating the refined representations obtained through iterative inference. This approach has the advantage of maintaining stability during learning and prevents the accumulation of errors in deeper layers, which is crucial for scaling PC networks, but also the disadvantage of storing information in memory, which is then used for the weight update only, making it less bio-plausible than the original formulation. A similar energy function has been used in a different way in a previous work, where the authors used it to guide the update of the neural activities instead of the weight updates (Whittington & Bogacz, 2017).
4.1.1 Experiments
Figure 3: Test accuracies of various algorithms on the CIFAR10 dataset, evaluated on models of different depths. From the second plot onward, each pair of bars compares the performance of the algorithm with and without center nudging (CN).
Our hypothesis is that the proposed methods improve the performance of PC and iPC on deep models. To this end, we perform experiments on VGG-like models (Simonyan & Zisserman, 2014) — convolutional models followed by feedforward layers — on the CIFAR10 dataset, where we observe their test accuracies as a function of their depth. We again use backprop, PC, and iPC as baselines, and report the results computed with and without centered nudging in Figure 3. All the details needed to reproduce the experiments, architectures, and hyperparameters used can be found in the supplementary material.
Results.
The barplots show that PC and iPC with and without center nudging significantly drop in test accuracy when the depth of the model is increased. In contrast, our proposed methods avoid the accuracy degradation as model depth increases. Particularly, iPC with spiking precision is the method that exhibits performance comparable to BP across VGG5/7/10 models, despite being completely local in space and time. We will later see that this is consistent with the more complex experiments, where this method is still the one achieving the best performance overall. In models without forward update, consistent with previous findings, using center nudging enhances algorithm accuracy, with more pronounced effects as the number of model layers increases in most cases. However, we found that once forward update is added, center nudging ceases to yield performance benefits.
Table 1: Performance of different precision schedules on VGG10/CIFAR-10.
Method Test Accuracy (%)
Fixed + FU $87.33\pm 0.14$
Lin Decay + FU $88.35\pm 0.12$
Exp Decay + FU $89.43\pm 0.18$
Spiking + FU $92.07\pm 0.10$
Linear and exponential decays.
To address the problem of energy imbalances, we used spikes that help propagate the energy to the first layers. However, alternative options would have been to attenuate the energy accumulation in the last layers by gradually reducing their precision over time. We therefore evaluate two variants where precisions decay either linearly or exponentially with the number of time steps, using the same setup of the experiment above, that is, a VGG10 on CIFAR10. The results, reported in Tab. 1, show that such decays provide improvements over the baseline, confirming the benefit of dynamically modulating precisions. However, spiking precisions consistently outperform both decay schedules, underscoring that actively boosting error propagation is more effective than passively damping energy imbalances. We refer to the supplementary material for a more detailed description.
4.2 Structural Contributions
Residual Connections.
Previously, we showed that our proposed methods overcome the depth limitation when training VGG-style models with $10$ layers. However, this improvement does not extend to ResNet10, where performance still drops catastrophically. We conjecture that this degradation arises from a mismatch in how energy propagates through residual connections: the skip path delivers energy to higher layers earlier than the main feedforward path. Concretely, consider the residual block in Fig. 4(a): when the neurons $\mathbf{x}^{l+2}$ receive energy at time $t$, in the next step they propagate it both to $\mathbf{x}^{l+1}$ (through the feedforward connection) and to $\mathbf{x}^{l-1}$ (through the skip connection). As a result, higher-level neurons begin updating before the main stream of energy arrives.
Having two streams of energy that reach the same levels in different time steps contrasts with both the philosophy behind the spiking updates, that are supposed to boost the neural activities the first time this is reached by the energy, and iPC, that updates some weight parameters before the information that goes through the model has reached them. Furthermore, it has been shown that it makes the feedback signal diverge from the one of backprop Salvatori et al. (2022). To address this temporal mismatch, we introduce auxiliary families of neurons along each residual connection, one for every skipped layer, as shown in Fig. 4(a). These auxiliary units act as buffers that delay the propagation of energy through the skip path, ensuring that energy arriving via the residual connection reaches the start of the block synchronously with the energy propagated through the main path.
Figure 4: (a): A sketch of a residual block (left) and our proposed variation (right) with auxiliary neural activities, that prevent the error signal to travel from $\mathbf{x}^{l+2}$ to $\mathbf{x}^{l-1}$ in one timestep; (b) A barplot showing the gap in test accuracy on the CIFAR10 dataset between models with and without added neural activities on a ResNet18. Both plots refer to the best test accuracies reached by PC and iPC with the novel methods presented above. (c) Shows the test accuracy between models with the standard formulation of BN, without BN, and our proposed BF.
@@ 4.2 Structural Contributions
where $\gamma$ and $\beta$ are learnable parameters, and $\mu_{B}$ and $\sigma^{2}_{B}$ are the mean and variance of the minibatch $B$. At test time, it uses statistics $\mu_{r}$ and $\sigma^{2}_{r}$ learned through exponential moving averages. However, when applied directly to PCNs, BatchNorm fails to yield similar improvements. We hypothesize that this failure results from the iterative inference phase, where processing the same batch multiple times leads to a possible overfitting of the layer statistics. To address this issue, we propose BatchNorm Freezing (BF), a modification that freezes the states of the BatchNorm state during the inference phase, and updates running statistics exclusively during the learning phase while still using batch statistics for normalization during inference iterations.
Experiments.
With the goal of validating the conjectures outlined above, we conducted two sets of experiments on the CIFAR-10 dataset: First, we trained a ResNet10 and a ResNet18 both with and without auxiliary neurons on the residual connections. In this setting, we used the two best-performing training schemes identified earlier, iPC with spiking updates, and PC with spiking plus forward updates. Second, we revisited the VGG10 architecture and compared different forms of batch normalization, including our newly proposed BF variant. The results of both experiments are reported in Fig. 4(b) and (c), respectively.
Results.
For ResNet10, the results indicate that introducing auxiliary neurons resolves the skip-connection issue, enabling both algorithms to match the performance of the VGG10 baseline reported previously. This result is even starker in ResNet18, where the presence of multiple skip connections causes a much larger degradation of performance. Again, the addition of auxiliary neurons completely solves the problem, as the proposed models reach almost $92\%$ test accuracy, compared to a maximum of $64\%$. A similar result is obtained in the experiments with BF, where we show that normalization not only maintains convergence but also improves accuracy, indicating that stabilizing activation distributions plays a crucial role in supporting energy-based optimization within predictive coding architectures. In the next section we will show that such results, together with the ones obtained at the beginning of the section, will also extend to deeper models and more complex datasets.
@@ 5 Experiments
Here we test our proposed methods combined, and show that we are able to reach performance comparable to those of BP when trained on models with the same complexity. To provide a comprehensive evaluation, we test them on CIFAR-10/100 (Krizhevsky et al., 2009) and TinyImageNet (Le & Yang, 2015), a scaled down version of ImageNet with $200$ classes. As architectures, we use VGG-like models and ResNet architectures (He et al., 2016). Similarly to the setups of the aforementioned works, in this work we only consider models with a single feedforward layer after the convolutional layers. In all cases, we perform a large number of hyperparameter searches, and report the best test accuracy obtained with early stopping, averaged over $5$ runs. We have run the experiments with and without BN/BF, and reported the best one obtained. Details on the architecture used, information needed to reproduce the results, study on hyperparameter importance, are in Appendix B.
@@
Table 2: Test accuracies of different algorithms across datasets and architectures.
Dataset Algorithm VGG5 VGG7 VGG10 ResNet10 ResNet18
CIFAR10 BP $\mathbf{90.01^{\pm 0.15}}$ $\mathbf{91.32^{\pm 0.14}}$ $92.68^{\pm 0.10}$ $\mathbf{92.93^{\pm 0.11}}$ $\mathbf{93.13^{\pm 0.16}}$
PC $87.98^{\pm 0.11}$ $84.62^{\pm 0.10}$ $76.22^{\pm 0.43}$ $69.85^{\pm 2.12}$ $15.63^{\pm 7.22}$
iPC $86.01^{\pm 0.10}$ $80.15^{\pm 0.18}$ $63.83^{\pm 0.33}$ $62.34^{\pm 0.27}$ $21.90^{\pm 1.51}$
iPC + Spiking Updates $89.73^{\pm 0.06}$ $91.12^{\pm 0.08}$ $93.03^{\pm 0.18}$ $92.39^{\pm 0.04}$ $91.96^{\pm 0.07}$
PC + Spiking + Forward $89.45^{\pm 0.18}$ $90.89^{\pm 0.04}$ $\mathbf{93.27^{\pm 0.10}}$ $92.47^{\pm 0.01}$ $91.93^{\pm 0.14}$
CIFAR100 (Top-1) BP $\mathbf{67.39^{\pm 0.25}}$ $67.67^{\pm 0.11}$ $71.25^{\pm 0.21}$ $71.21^{\pm 0.09}$ $\mathbf{71.69^{\pm 0.21}}$
PC $61.84^{\pm 0.18}$ $56.80^{\pm 0.14}$ $50.76^{\pm 0.37}$ $41.51^{\pm 0.32}$ $1.59^{\pm 0.02}$
iPC $56.07^{\pm 0.16}$ $43.99^{\pm 0.30}$ $31.99^{\pm 0.17}$ $22.91^{\pm 0.23}$ $1.56^{\pm 0.24}$
iPC + Spiking Updates $66.91^{\pm 0.12}$ $67.10^{\pm 0.12}$ $69.84^{\pm 0.17}$ $70.02^{\pm 0.24}$ $70.38^{\pm 0.20}$
PC + Spiking + Forward $67.16^{\pm 0.16}$ $\mathbf{67.71^{\pm 0.10}}$ $\mathbf{72.02^{\pm 0.12}}$ $\mathbf{71.30^{\pm 0.21}}$ $70.90^{\pm 0.18}$
CIFAR100 (Top-5) BP $89.56^{\pm 0.08}$ $\mathbf{90.05^{\pm 0.13}}$ $\mathbf{92.10^{\pm 0.12}}$ $89.49^{\pm 0.13}$ $89.43^{\pm 0.14}$
PC $86.53^{\pm 0.15}$ $83.00^{\pm 0.09}$ $78.68^{\pm 0.27}$ $70.95^{\pm 0.35}$ $5.89^{\pm 0.12}$
iPC $78.91^{\pm 0.23}$ $73.23^{\pm 0.30}$ $61.17^{\pm 0.31}$ $46.41^{\pm 0.31}$ $6.33^{\pm 0.26}$
iPC + Spiking Updates $89.47^{\pm 0.02}$ $9.42^{\pm 0.06} 88$$75^{\pm 0.29} 88.$$8.90^{\pm 0.21}$ $\mathbf{90.74^{\pm 0.10}}$
PC + Spiking + Forward $\mathbf{89.57^{\pm 0.09}}$ $89.62^{\pm 0.18}$ $\mathbf{92.10^{\pm 0.10}}$ $\mathbf{89.56}^{\pm 0.1}$ $90.47^{\pm 0.11}$
TinyImageNet (Top-1) BP $47.81^{\pm 0.12}$ $50.13^{\pm 0.06}$ $53.61^{\pm 0.12}$ $53.02^{\pm 0.20}$ $\mathbf{58.18^{\pm 0.12}}$
PC $41.29^{\pm 0.20}$ $41.15^{\pm 0.14}$ $31.87^{\pm 0.03}$ $13.59^{\pm 0.12}$ $0.84^{\pm 0.02}$
iPC $29.94^{\pm 0.47}$ $9.76^{\pm 0.15} 11$$41^{\pm 0.23} 11.$$1.66^{\pm 0.39}$ $1.44^{\pm 0.05}$
iPC + Spiking Updates $48.13^{\pm 0.10}$ $48.75^{\pm 0.12}$ $52.40^{\pm 0.20}$ $\mathbf{54.94^{\pm 0.30}}$ $57.83^{\pm 0.21}$
PC + Spiking + Forward $\mathbf{49.35^{\pm 0.09}}$ $\mathbf{50.64^{\pm 0.12}}$ $\mathbf{55.31^{\pm 0.25}}$ $53.25^{\pm 0.27}$ $54.24^{\pm 0.63}$
TinyImageNet (Top-5) BP $72.15^{\pm 0.10}$ $3.94^{\pm 0.10} 77$$7.11^{\pm 0.10}$ $72.90^{\pm 0.16}$ $\mathbf{79.94^{\pm 0.06}}$
PC $6.68^{\pm 0.09} 66$$6.25^{\pm 0.11}$ $58.14^{\pm 0.04}$ $37.99^{\pm 0.08}$ $5.34^{\pm 0.01}$
iPC $54.73^{\pm 0.52}$ $40.36^{\pm 0.22}$ $30.42^{\pm 0.36}$ $26.51^{\pm 0.71}$ $11.58^{\pm 0.21}$
iPC + Spiking Updates $72.71^{\pm 0.12}$ $73.39^{\pm 0.10}$ $76.76^{\pm 0.15}$ $\mathbf{78.88^{\pm 0.32}}$ $77.55^{\pm 0.12}$
PC + Spiking + Forward $\mathbf{73.46^{\pm 0.09}}$ $\mathbf{75.63^{\pm 0.08}}$ $\mathbf{79.30^{\pm 0.17}}$ $77.72^{\pm 0.23}$ $74.70^{\pm 0.47}$
Results.